OpenStudy (anonymous):

Which expression can be used to solve the problem below? To cater a luncheon a hotel charger $50 per hour for use of a dining room plus $24.50 per person. What is the total cost for a 2-hour luncheon for 45 people?

OpenStudy (anonymous):

Please explain.

OpenStudy (anonymous):

$50 per hour (2 hours) $24.50 per person (45 people)

OpenStudy (anonymous):

Its asking what info is needed to solve the problem

OpenStudy (anonymous):

Oh never mind, sorry

OpenStudy (anonymous):

its asking what equation could you use to find the total cost

Parth (parthkohli):

Take x as the cost of dining room per hour(x = 50) Take y as the cost of the individuals(y = 24.50) 2x + 45y = the cost =100+1102.5 = $1202.5

OpenStudy (anonymous):

(24.50 *X) + (50 * Y) = T where T is the total cost where X is the number of people where Y is the number of hours

OpenStudy (anonymous):

@ParthKohli It's better to use for variables things that will change. The rate will rarely change while it's easy to change the number of hours and the number of people.

Parth (parthkohli):

@Romero I first put the equation and then did what the question required and solved it then.

OpenStudy (anonymous):

@ParthKohli In real life situations that equation will be wrong to use.

OpenStudy (anonymous):

would that be 2 X50 + 24.50*45?

Parth (parthkohli):

Yes @happyhorse

OpenStudy (anonymous):

oK, THANKS, PATHKOHLI AND ROMERO.

Parth (parthkohli):

I put the equation with variables generally. Then I substituted. What's the problem?

OpenStudy (anonymous):

Im going to fan you both. thanks.

OpenStudy (anonymous):

@ParthKohli Because it's the wrong way to do it. It's just fact.

Parth (parthkohli):

@happyhorse Thanks

OpenStudy (anonymous):

no prob, thannx for fanning me!